By how much does a 54-kg mountain climber stretch her 0.871-cm diameter nylon rope when she hangs 49 m below a rock outcropping? (For nylon, Y=1.35 × 109Pa.) 


1
Expert's answer
2021-03-26T11:41:18-0400

We can find the Young's modulus as follows:


"E=\dfrac{FL_0}{A\Delta L},""\Delta L=\dfrac{FL_0}{AE}=\dfrac{mgL_0}{\dfrac{\pi d^2}{4}E},""\Delta L=\dfrac{54\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot49\ m}{\dfrac{\pi (0.871\cdot10^{-2}\ m)^2}{4}\cdot1.35\cdot10^9\ Pa}=0.32\ m."